# Archilles’ Heel in Semi-open LLMs: Hiding Bottom against Recovery Attacks

### v1

@@ 4.1 Resilience Transition Layer in Infinitely Deep Transformers
Model Overview. Let us revisit our large language model composed of $L$ layers, denoted as $f(\mathbf{X};\bm{\theta})=\varphi_{L}\circ...\circ\varphi_{1}(\mathbf{X})$. Recall that each row of the feature matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$ represents a $d$-dimensional vector for an input token. We treat each layer $\varphi_{i}$ as a transformer layer, where each layer processes an $n\times d$ dimensional matrix as input and outputs another $n\times d$ matrix. Thus, the model $f$ outputs a matrix of $n$ rows and $d$ columns, indicating that the large language model outputs a feature vector for each token. Moreover, we assume that each layer contains a normalized residual self-attention function, defined as
@@
$\varphi_{i}\left(\mathbf{X};K_{i},Q_{i}\right)=\mathbf{X}+\operatorname{softmax}\left(\frac{\mathbf{X}Q_{i}(\mathbf{X}K_{i})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)\mathbf{X},$ (2)
@@
where $Q_{i}\in\mathbb{R}^{d\times d_{Q}}$ and $K_{i}\in\mathbb{R}^{d\times d_{Q}}$ are projection parameter matrices for the $Q$ and $K$ matrices in the transformer, respectively. Additionally, $\sqrt{d_{Q}}$ and the matrix norm $\|\mathbf{X}\|$ denote normalization factors provided by the normalization layer. We consider the strategy of concealing the $\alpha L$-th layer with $\alpha\in[0,1]$ and $\alpha L\in\mathbb{N}$ while keeping other layers public. After the semi-open model recovery, we assume the parameters of the recovered model in the public layers are identical to the victim model, while those in the proprietary layer deviate. Let $\hat{K}_{\alpha L}$ and $\hat{Q}_{\alpha L}$ denote the recovered weight matrix of the proprietary layer, i.e., ${\bm{\theta}}_{\text{FT}}(\{\alpha L\})=\{(K_{1},Q_{1}),...,(\hat{K}_{\alpha L},\hat{Q}_{\alpha L}),...,({K}_{L},{Q}_{L})\}$. Let $\hat{\varphi}_{\alpha L}$ denote the function of the recovered proprietary layer, i.e., the $\alpha L-$th layer, in the recovered model. In this subsection, we consider the normalized output of an infinitely deep model whose $\alpha L$-th layer is closed-sourced and subjected to the attack. The output of the recovered model is
@@
$\hat{f}_{\infty}(\mathbf{X})=\lim_{L\rightarrow\infty}\frac{\varphi_{L}\circ...\varphi_{\alpha L+1}\circ\hat{\varphi}_{\alpha L}\circ\varphi_{\alpha L-1}\circ...\circ\varphi_{1}(\mathbf{X})}{\|\varphi_{L}\circ...\varphi_{\alpha L+1}\circ\hat{\varphi}_{\alpha L}\circ\varphi_{\alpha L-1}\circ...\circ\varphi_{1}(\mathbf{X})\|_{F}},$
@@
where $\|\cdot\|_{F}$ denotes the Frobenius norm of a given matrix. We consider this infinitely deep network as our ideal model because, in real-world settings, most large-scale models are sufficiently deep. Next, we present the following theorem to illustrate the existence of a critical value $\alpha^{*}$ such that if $\alpha<\alpha^{*}$, the recovered LLM outputs identical feature vectors for all tokens. Conversely, if $\alpha>\alpha^{*}$, the output feature vectors may vary across tokens.
@@
(1) Let $\alpha<\alpha^{\star}$ and $\{K_{i},Q_{i}\}_{i\geq 1}$ be any parameter matrix sequence in the victim model. Let $\hat{K}_{\alpha L}$ and $\hat{Q}_{\alpha L}$ be the recovered parameter matrices drawn from a continuous distribution supported on $\mathbb{R}^{n\times d}$. With probability one, for any input $\mathbf{X}\in\mathcal{X}$, the row vectors in the matrix $\hat{f}_{\infty}(\mathbf{X})$ are identical.
@@
(2) Let $\alpha>\alpha^{\star}$. There exists a victim model with parameter matrix sequence $\{K_{i},Q_{i}\}_{i\geq 1}$ such that for any recovered parameter matrices $\hat{K}_{\alpha L}$ and $\hat{Q}_{\alpha L}$, the row vectors in the matrix $\hat{f}_{\infty}(\mathbf{X})$ are not entirely the same for some input feature matrix $\mathbf{X}\in\mathcal{X}$.
@@
Remark 1: The proof is provided in Appendix A. This theorem demonstrates that if the recovered parameters of the bottom layers (i.e., $\alpha<\alpha^{*}$) are obtained through a randomized algorithm, such as stochastic gradient descent, with a continuous distribution supported on $\mathbb{R}^{n\times d}$, the recovery will certainly fail, as it will produce the same feature vector for every token. In contrast, keeping the later layers closed-sourced (i.e., $\alpha>\alpha^{*}$) does not maintain this property, indicating that it is more effective to closed-source the bottom layers before the transition layer, rather than the later ones.
Remark 2: The theorem relies on the assumption that the distribution is defined over a countable domain, $\mathcal{X}\times\mathcal{Y}$, typically satisfied by inputs such as sentences or images. We show in the proof that for each input matrix $\mathbf{X}\in\mathcal{X}$, there are two zero-measure sets $\mathcal{K}(\mathbf{X})$ and $\mathcal{Q}(\mathbf{X})$ such that the recovered matrices must avoid to satisfy the theorem. Hence, the countable unions $\mathcal{K}=\bigcup_{\mathbf{X}\in\mathcal{X}}\mathcal{K}(\mathbf{X})$ and $\mathcal{Q}=\bigcup_{\mathbf{X}\in\mathcal{X}}\mathcal{Q}(\mathbf{X})$ are also zero-measure sets, ensuring that when recovered matrices do not belong to these sets, the conditions in the theorem are met for any input matrix $\mathbf{X}$ in the input space.
Theorem 1 shows that hiding bottom layers offers greater resilience against model recovery attacks, indicating that closed-sourcing layers starting from the first layer may be an effective approach. In the next subsection, we present a closed-sourcing approach to identify the minimal set of hidden layers.
4.2 SCARA: Selective Closed-sourcing Approach Against Recovery Attack
@@
We propose a method to approximately find the smallest bottom layer index set $I$ that satisfies $R(I)\leq(1+\varepsilon)R([L])$. A simple approach is to start with $I_{l}=\{1,\ldots,l\}$ for each $l$ beginning from 1, then evaluate the recovery ratio $R(I_{l})$ after the attack, and identify the smallest $l$ that meets the inequality. This extensive fine-tuning process is time-consuming, prompting the critical question: Can we create a fine-tuning-free metric that predicts LLM performance under semi-open model recovery attacks? Hence, our goal is to establish a metric directly correlated with the recovery ratio.
@@
In the recovery ratio $R(I)$, each $I$ has the same denominator, so our focus is on a metric related to the numerator, specifically $\mathbb{E}[s(f(\mathbf{X};\bm{\theta}_{\text{FT}}(I,\mathcal{D})),Y)]$, which measures the average performance score of the recovered model. This average performance score generally inversely correlates with the average testing loss $\mathbb{E}[\ell(f(\mathbf{X};\bm{\theta}_{\text{FT}}(I,\mathcal{D})),Y)]$. Therefore, our goal becomes finding the $l$ such that
$\mathbb{E}[\ell(f(\mathbf{X};\bm{\theta}_{\text{FT}}(\{1,...,l\},\mathcal{D})),Y)]\geq(1-\varepsilon)\mathbb{E}[\ell(f(\mathbf{X};\bm{\theta}_{\text{FT}}([L],\mathcal{D})),Y)].$
@@
However, calculating both sides of this inequality requires knowing the recovered parameters from the fine-tuning process. To bypass this, we aim for an approximate solution. The recovered parameters are generated through gradient descent, starting from the initial parameters $\bm{\theta}_{0}(I)$, with the hidden layers being randomly initialized. Using the Taylor Expansion, we find
$\mathbb{E}\left[\ell\left(f(\mathbf{X};\bm{\theta}_{\text{FT}}(I,\mathcal{D})),Y\right)\right]=\mathbb{E}\left[\ell\left(f(\mathbf{X};\bm{\theta}_{0}(I)),Y\right)\right]+\mathcal{O}(\|\bm{\theta}_{\text{FT}}(I,\mathcal{D})-\bm{\theta}_{0}(I)\|_{2}).$
@@
Previous research (Choi et al., 2024; Bailly et al., 2022) suggests the difference $\|\bm{\theta}_{\text{FT}}(I,\mathcal{D})-\bm{\theta}_{0}(I)\|_{2}$ is minor for large networks compared to the dataset size $|\mathcal{D}|$. In models like a single-layer ReLU network (Anthony et al., 1999; Zou et al., 2020), the difference $\|\bm{\theta}_{\text{FT}}(I,\mathcal{D})-\bm{\theta}_{0}(I)\|_{2}$ is of order $\mathcal{O}\left(\frac{|\mathcal{D}|}{\sqrt{N}}\right)$ (Jacot et al., 2018; Wei et al., 2019), where $N$, the number of model parameters, which is much larger than the dataset size in LLMs (Dubey et al., 2024; Liu et al., 2024). Hence, the first term that does not require fine-tuning dominates, suggesting it as a viable metric for predicting the recovery ratio. Thus, we define the first term as “Recovery Difficulty” (RD$(I)$) with the expression:
$\text{RD}(I)=\mathbb{E}_{\mathbf{X},Y,\bm{\theta_{0}}(I)}\left[\ell\left(f(\mathbf{X};\bm{\theta}_{0}(I),Y\right)\right].$
@@
This score, which can be estimated using a sample average, represents the recovered performance of the model when specific layers $I$ are closed-sourced. A higher RD$(I)$ suggests worse recovery performance, indicating a lower recovery ratio $R(I)$. Therefore, our SCARA operates in the following way. SCARA begins by sampling evaluation data from the underlying distribution and then computes RD$(I_{l})$ for each set of closed-sourced layers $I_{l}=\{1,...,l\}$ for $l=1,...,L$. SCARA stops at the smallest $l^{*}$ that satisfies RD$(I_{l^{*}})\geq(1-\varepsilon)\text{RD}([L])$.
@@ 5.1 Experimental Settings
Models. We consider five open-source, decoder-only structured LLMs with various architectures. Specifically, we select Llama2-70B-chat, Llama2-7B-chat (Touvron et al., 2023), Mistral-7B-v0.1 (Jiang et al., 2023a), Phi-2 (Abdin et al., 2024), and Phi-1.5 (Li et al., 2023). We designate these pre-trained models as the base models for customization and victims in semi-open model recovery attacks. Details of each model is provided in Appendix B.1.
@@
Attack Methods. We recover models produced by different closed-source approaches using three attack methods: FT-all, FT-closed, and SEM (Stealing Embedding Model) (Tamber et al., 2024). FT-all finetunes the entire model, whereas FT-closed fine-tunes the closed-sourced set only. We also consider SEM, a typical attack that trains a substitute embedding module with data from the closed-sourced set. Following (He et al., 2021), a diverse attack set is required for full recovery. Therefore, we merge data evenly for two general datasets, MMLU benchmark (Hendrycks et al., 2021) and Alpaca 52k (Wang et al., 2022), resulting in a 51k combined set. Details are in Appendix B.2
Baselines. We compare SCARA with the other two baselines: SAP-DP and the fully-closed (Eiras et al., 2024) approach. The SAP (Shen et al., 2023) framework keeps the first six decoder layers open and the rest closed-source. SAP-DP extends SAP by adding Laplace noise to the model outputs, a common strategy for model protection (Lee et al., 2018). The fully-closed approach represents the extreme, where all layers are closed-sourced. Further details can be found in Appendix B.3
Implementation Details of SCARA. We apply the SCARA algorithm to identify the smallest closed-sourced set $I$ such that $R(I)\leq(1+\varepsilon)R([L])$. To calculate the recovery difficulty (RD), we use a 1.5k evaluation set sourced from the same datasets as the attack set. We explore $\varepsilon$ values from 0.05 to 1 and discover that setting $\varepsilon$ to 0.05 yields optimal performance on the evaluation set. We randomly initialize the hidden layers and average the RD across three seeds to select the smallest layer set. For more details on the dataset and the SCARA sensitivity on $\varepsilon$, see Appendix B.4 and Section 5.3.
Evaluation Benchmarks We assess customizability on six downstream tasks: Code (Zheng et al., 2024b), Math (Yue et al., 2023), Medical (Zhang et al., 2023), Finance (Wang et al., 2023b), Law (Guha et al., 2024), and Alignment (Meng et al., 2024). To fully evaluate recovered functionalities, we focus on six capabilities domains following Llama2 report (Touvron et al., 2023). Specifically, we assess the recovered model across sixteen benchmarks grouped into (1) Commonsense Reasoning (Rsn.); (2) Reading Comprehension (Read.); (3) World Knowledge (Knl.); (4) Code; (5) Math; and (6) General Ability (Gen.). More details can be found in Appendix B.5.
Metrics. We measure model customizability through its improvements on benchmarks. For resilience, we calculate the “Average Recovery Ratio” (ARR) by averaging the recovery ratios across benchmarks. A lower ARR indicates higher resilience offered by the closed-sourced set. Additionally, we define $\Delta\textbf{ARR}(I)=\text{ARR}(I)-\text{ARR}([L])$ to compare the resilience between closed-sourcing set $I$ and the fully-closed approach. A smaller $\Delta\textbf{ARR}$ suggests similar resilience to the fully-closed model.
Figure 4: Customization performance of models closed-sourced by SCARA on six downstream tasks.

### v2

@@ 4.1 Security Transition in Deep Transformers
Model Overview. In this subsection, we consider a deep transformer $f$ with $L$ layers, expressed as $f(\mathbf{X};\bm{\theta})=\varphi_{L}\circ\dots\circ\varphi_{1}(\mathbf{X})$. The input feature matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$ consists of $n$ rows, each representing a $d$-dimensional token vector. Each layer $\varphi_{i}$ is a transformer that incorporates a normalized residual self-attention mechanism, defined as:
@@
$\varphi_{i}(\mathbf{X};K_{i},Q_{i})=\mathbf{X}+\operatorname{softmax}\left(\frac{\mathbf{X}Q_{i}(\mathbf{X}K_{i})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)\mathbf{X}.$
@@
Here, $Q_{i}\in\mathbb{R}^{d\times d_{Q}}$ and $K_{i}\in\mathbb{R}^{d\times d_{Q}}$ are projection matrices for the query and key components, respectively. The terms $\sqrt{d_{Q}}$ and $\|\mathbf{X}\|$ serve as normalization factors, ensuring stable computations within the attention mechanism. We consider the semi-open deployment of securing the $\alpha L$-th layer with $\alpha\in[0,1]$ and $\alpha L\in\mathbb{N}$ while keeping other layers unsecured. After the distillation attack, we assume the parameters of the distilled model in the unsecured layers are identical to the victim model, while those in the secured layer deviate. Let $\hat{K}_{\alpha L}$ and $\hat{Q}_{\alpha L}$ denote the distilled weight matrix of the proprietary layer, i.e., ${\bm{\theta}}_{\text{dist}}(\{\alpha L\})=\{(K_{1},Q_{1}),...,(\hat{K}_{\alpha L},\hat{Q}_{\alpha L}),...,({K}_{L},{Q}_{L})\}$. Let $\hat{\varphi}_{\alpha L}$ denote the function of the distilled proprietary layer, i.e., the $\alpha L-$th layer, in the distilled model. In this subsection, we consider the normalized output of an infinitely deep model whose $\alpha L$-th layer is hidden and subjected to the attack. The output of the distilled model is
@@
$\hat{f}_{\infty}(\mathbf{X})=\lim_{L\rightarrow\infty}\frac{f(\mathbf{X};{\bm{\theta}}_{\text{dist}}(\{\alpha L\}))}{\|f(\mathbf{X};{\bm{\theta}}_{\text{dist}}(\{\alpha L\}))\|_{F}},$
@@
where $\|\cdot\|_{F}$ denotes the Frobenius norm. We consider an infinitely deep network as the ideal model, reflecting the sufficient depth of most large-scale models in practice. The following theorem establishes the existence of a critical value $\alpha^{*}$ such that if $\alpha<\alpha^{*}$, the output matrix of the distilled LLM has rank one. Conversely, if $\alpha>\alpha^{*}$, the output matrix has rank strictly greater than one.
@@
(1) If $\alpha<\alpha^{*}$ and $\{K_{i},Q_{i}\}_{i\geq 1}$ are parameter matrices of the victim model, with $\hat{K}_{\alpha L}$ and $\hat{Q}_{\alpha L}$ as distilled parameters drawn from a continuous distribution on $\mathbb{R}^{n\times d}$, the matrix $\hat{f}_{\infty}(\mathbf{X})$ exists and almost surely has rank one for all $\mathbf{X}$.
@@
(2) If $\alpha>\alpha^{*}$, there exists a victim model with parameter sequence $\{K_{i},Q_{i}\}_{i\geq 1}$ such that for any distilled parameters $\hat{K}_{\alpha L}$ and $\hat{Q}_{\alpha L}$, the matrix $\hat{f}_{\infty}(\mathbf{X})$ exists but has rank greater than one for some $\mathbf{X}$.
@@
Remark 1: The proof is provided in Appendix A. This theorem demonstrates that if the distilled parameters of the bottom layers (i.e., $\alpha<\alpha^{*}$) are obtained through a randomized algorithm, such as stochastic gradient descent, with a continuous distribution supported on $\mathbb{R}^{n\times d}$, the distillation will certainly fail, as the feature matrix degenerate. In contrast, keeping the later layers secured (i.e., $\alpha>\alpha^{*}$) does not maintain this property, indicating that it is more effective to secure the bottom layers before the transition layer, rather than the later ones.
Remark 2: The existence of $\hat{f}_{\infty}(\mathbf{X})$ is a non-trivial result. While the mapping $\varphi_{i}$ admits a fixed point at $\mathbf{X}=\mathbf{0}_{n\times d}$, the convergence of the iterative process governed by $\varphi_{i}$ cannot be guaranteed using the contraction mapping theorem, as $\varphi_{i}$ does not satisfy the contraction property for any pair $(Q_{i},K_{i})$. This complexity becomes particularly evident in the special case where $n=1$ and $\mathbf{X}$ is a column vector. Here, the output of $\varphi_{i}$ satisfies the relation $\langle\mathbf{1}_{d},\varphi_{i}(\mathbf{X};K_{i},Q_{i})\rangle=2\langle\mathbf{1}_{d},\mathbf{X}\rangle$, implying that the iteration diverges unless $\mathbf{X}$ is orthogonal to $\mathbf{1}_{d}$. However, the divergence is not arbitrary; rather, the theorem reveals that it occurs in a fixed, well-defined direction. This insight ensures the existence of a normalized output, which remains stable and meaningful despite the lack of strict convergence.
Remark 3: The existence of $\alpha^{*}\in(0,1)$ is also a non-trivial statement, as $\alpha^{*}$ could potentially be zero, which would imply the absence of a critical layer such that securing layers prior to it guarantees the failure of the recovered model’s functionality. The primary challenge lies in demonstrating that perturbations to the earlier layers result in rank-one outputs, a property that does not universally hold for arbitrary perturbations. To address this, we establish an alternative result: given an input matrix $\mathbf{X}$, rank-one outputs can be guaranteed if the perturbation matrices $K_{i}$ and $Q_{i}$ are chosen to avoid specific zero-measure sets, denoted as $\mathcal{K}(\mathbf{X})$ and $\mathcal{Q}(\mathbf{X})$, respectively. Assuming a countable domain $\mathcal{X}\times\mathcal{Y}$, which is typical for structured inputs such as sentences or images, it follows that the perturbation matrices to be avoided belong to the countable union of these sets, defined as $\mathcal{K}=\bigcup_{\mathbf{X}\in\mathcal{X}}\mathcal{K}(\mathbf{X})$ and $\mathcal{Q}=\bigcup_{\mathbf{X}\in\mathcal{X}}\mathcal{Q}(\mathbf{X})$. Since this union remains a zero-measure set, avoiding these specific sets ensures that the conditions of the theorem are satisfied for any input matrix $\mathbf{X}$.
Theorem 1 shows that protecting bottom layers enhances resilience against model distillation attacks, indicating that securing layers starting from the first may be effective. In the next subsection, we propose SOLID, a preliminary LLM deployment framework solution that balance model protection and customization.
4.2 SOLID: Semi-Open Local Infrastructure Deployment Framework
@@
We propose a method to approximately find the smallest bottom layer index set $I$ that satisfies $R(I)\leq(1+\varepsilon)R([L])$ for any small $\varepsilon>0$. A simple approach is to start with $I_{l}=\{1,\ldots,l\}$ for each $l$ beginning from 1, then evaluate the distillation ratio $R(I_{l})$ after the attack, and identify the smallest $l$ that meets the inequality. This extensive fine-tuning process is time-consuming, prompting the critical question: Can we create a fine-tuning-free metric that predicts LLM performance under model distillation attacks? Hence, our goal is to establish a metric directly correlated with the distillation ratio.
@@
In the distillation ratio $R(I)$, each $I$ has the same denominator, so our focus is on a metric related to the numerator, specifically $\mathbb{E}[s(f(\mathbf{X};\bm{\theta}_{\text{FT}}(I,\mathcal{D})),Y)]$, which measures the average performance score of the distilled model. This average performance score generally inversely correlates with the average testing loss with the expression $L(\bm{\theta})\triangleq\mathbb{E}_{\mathbf{X}\times Y}[\ell(f(\mathbf{X};\bm{\theta}),Y)]$, where $\ell$ denotes the cross-entropy loss employed by LLM. Hence, we aim at finding the smallest $I$ such that
$L({\bm{\theta}}_{\text{dist}}(I,\mathcal{D}))\geq(1-\varepsilon)L({\bm{\theta}}_{\text{dist}}([L],\mathcal{D})).$
@@
However, calculating both sides of this inequality requires knowing the distilled parameters from the fine-tuning process. To bypass this, we aim for an approximate solution. The distilled parameters are generated through gradient descent, starting from the initial parameters $\bm{\theta}_{0}(I)$, with the hidden layers being randomly initialized. Using the Taylor Expansion, we find
$L({\bm{\theta}}_{\text{dist}}(I,\mathcal{D}))$ $=L({\bm{\theta}}_{\text{0}}(I,\mathcal{D}))$
$\quad+\mathcal{O}(\mathbb{E}\|\bm{\theta}_{\text{dist}}(I,\mathcal{D})-\bm{\theta}_{0}(I)\|_{2}).$
@@
Previous research (Choi et al., 2024; Bailly et al., 2022) indicates that the difference $\|\bm{\theta}_{\text{dist}}(I,\mathcal{D})-\bm{\theta}_{0}(I)\|_{2}$ is minimal in large networks compared to the dataset size $|\mathcal{D}|$. In models such as single-layer ReLU networks (Anthony et al., 1999; Zou et al., 2020), this difference scales as $\mathcal{O}\left(\frac{|\mathcal{D}|}{\sqrt{N}}\right)$ (Jacot et al., 2018; Wei et al., 2019), where $N$, the number of model parameters, far exceeds the dataset size in large language models (LLMs) (Dubey et al., 2024; Liu et al., 2024). The first term, independent of fine-tuning, dominates and effectively predicts the distillation ratio. We refer to this term as the Distillation Difficulty (DD$(I)$), defined as
$\text{DD}(I)=\mathbb{E}[L(\bm{\theta}_{0}(I))].$
@@
This score, which can be estimated using a sample average, represents the distilled model performance of the model when specific layers $I$ are secured. A higher DD$(I)$ suggests better security performance, indicating a lower distillation ratio $R(I)$. Therefore, our SOLID operates in the following way. SOLID begins by sampling evaluation data targeting general capabilities from the underlying distribution, and then computes DD$(I_{l})$ for each set of secured layers $I_{l}=\{1,...,l\}$ for $l=1,...,L$. SOLID stops at the smallest $l^{*}$ that satisfies DD$(I_{l^{*}})\geq(1-\varepsilon)\text{DD}([L])$.
@@ 5 Experiments
In this section, we conduct experiments to validate the correctness of our positions. Specifically, we answer the following research questions:
  • RQ1. Can query-based distillation attacks distill the functionality of the entire model under the baseline deployment that secures the top layer?
  • RQ2. How do secured layer location and amount affect the security-customization trade-off?
  • RQ3. Does securing bottom layers (SOLID) offer a better balance between model theft risk and customization performance compared to baseline deployments?
@@ 5.1 Experimental Settings
We begin by introducing our experimental setups. Details can be found in Appendix B.
@@
Models. We consider five open-source, decoder-only structured LLMs with various architectures. Specifically, we select Llama2-70B-chat, Llama2-7B-chat (Touvron et al., 2023), Mistral-7B-v0.1 (Jiang et al., 2023), Phi-2 (Abdin et al., 2024), and Phi-1.5 (Li et al., 2023). We designate these pre-trained models as the base models for adaptation and victims in model distillation attacks.
@@
Attack Methods. We distill models produced by different protection approaches using three attack methods: FT-all, FT-closed and SEM. Following (He et al., 2021), a diverse attack set is required for full distillation. Therefore, we merge data evenly from two general datasets, MMLU (Hendrycks et al., 2021) and Alpaca 52k (Wang et al., 2022), resulting in a 51k combined set.
Baselines. We evaluate the effectiveness of SOLID by comparing it with three baselines: DarkneTZ (Mo et al., 2020), SAP-DP, and the fully-secured approach (Eiras et al., 2024). DarkneTZ protects only the final decoder layer while the SAP (Shen et al., 2023) framework exposes the first six decoder layers and secures the rest. SAP-DP extends SAP by adding Laplace noise to model outputs to enhance protection (Lee et al., 2018). The fully-secured approach represents the extreme, securing all layers for maximal security.
Implementation Details of SOLID. We apply the SOLID algorithm to identify the smallest secure set $I$ such that $R(I)\leq(1+\varepsilon)R([L])$. To calculate distillation difficulty (DD), we use cross-entropy loss and approximate the expectation over samples distributed on the general domain and randomly initialized secured parameters. This is done using a 1,500-sample evaluation set randomly sampled from the MMLU and Alpaca 52k, with secured parameters initialized via Xavier initialization and averaged over three random seeds (20, 42, 1234). In our experiments, we find that $\varepsilon=0.05$ yields optimal performance.